Two beads, one carrying charge +q and other carrying charge +4q, are separated by distance d that is much greater than the radius of each bead

Is ththere any location along the line between them where the electric field magnitude is zero?
1
Expert's answer
2021-09-04T18:21:26-0400

The electric field produced by the first bead at some distance "x" from it is given as follows:



"E_1 = k\dfrac{q}{x^2}"

where "k" is the constant and "x" is the distance from the first bead. The second bead produces the following field at the same point "x":




"E_2 = k\dfrac{4q}{(d-x)^2}"

These fields are directed in opposite direction (in the region between beads), thus, the resulting field is zero if



"E_1 = E_2"

Solving for "x", obtain:



"k\dfrac{q}{x^2} = k\dfrac{4q}{(d-x)^2}\\
\dfrac{1}{x^2} = \dfrac{4}{(d-x)^2}\\
(d - x)^2 = 4x^2\\
d^2 - 2dx + x^2 = 4x^2\\
3x^2 + 2dx - d^2 = 0\\
x = \dfrac{d + \sqrt{4d^2 +12d^2}}{6} = \dfrac{d + 4d}{6} = \dfrac{5}{6}d\\"



Answer. At distance "\dfrac{5}{6}d" from the first bead.